Suppose a​ ten-year, $ 1,000 bond with an 8.5 % coupon rate and semiannual coupons is trading for $ 1,035.23. a. What is the​ bond's yield to maturity​ (expressed as an APR with semiannual​ compounding)? b. If the​ bond's yield to maturity changes to 9.5 % ​APR, what will be the​ bond's price?
Answer #1

a.

the ​bond's yield to maturity using excel function

=RATE(nper,pmt,pv,fv)

=(RATE(10*2,1000*8.5%/2,-1035.23,1000))*2=7.98%

b.

the​ bond's price

=(1000*8.5%/2)*((1-(1+(9.5%/2))^(-2*10))/(9.5%/2))+1000/(1+(9.5%/2))^(2*10)

=936.35

the above will be answer..